You are offered a three year distribution contract for a technology product with a short market life. Your initial costs for the rights to sell the product are $2,500,000. Your first year net cash flow is estimated to be $3,400,000, after which it will decrease each year by 50% of the previous year’s net cash flow. You may assume that your interest rate for available investments is 9%.
Answer #1

Answer:

Initial Cost I=2,500,000

Net cash for first year CF1=3400000

CF2=3400000*(1-50%)=1700000

CF3=1700000*(1-50%)=850000

Interest rate r=9%

Net Present value of project NPV=-I+CF1/(1+r)^1+CF2/(1+r)^2+CF3/(1+r)^3

NPV=-2500000+3400000/(1+9%)+1700000/(1+9%)^2+850000/(1+9%)^3

NPV=$2706478

Since we are getting positive NPV so we should accept this project,